40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15672     Entry time: Tue Nov 10 17:46:06 2020
Author: gautam 
Type: Update 
Category: General 
Subject: IFO recovery 

Summary:

  1. Recovery was complicated by RFM failure on c1iscey - see Attachment #1. This is becoming uncomfortably frequent. As a result, the ETMY suspension wasn't being damped properly. Needed a reboot of c1iscey machine and a restart of the c1rfm model to fix.
  2. POX/POY locking was restored. Arm alignment was tuned using the dither alignment system.
  3. AS beam was centered on its CCD (I put a total of ND=1.0 filters back on the CCD). Note that the power in the AS beam is ~4x what it was given we have removed the in-vacuum pickoff to the OMC.
  4. Green beams were aligned to the arm cavities. See Attachment #2. Both green cameras were adjusted on the PSL table to have the beam be ~centered on them.
  5. ALS noise is far too high for locking, needs debugging. See Attachment #3.
  6. AS beam was aligned onto the AS55 photodiode. With the PRM aligned, the REFL beam was centerd on the various REFL photodiodes. The PRMI (resonant carrier) could be locked, see Attachment #4.

I want to test out an AS port WFS now that I have all the parts in hand - I guess the Michelson / PRMI will suffice until I make the ALS noise good again, and anyways, there is much assembly work to be done. Overnight, I'm repeating the suspension eigenmode measurement.

Attachment 1: RFMerrs.png  31 kB  | Hide | Hide all | Show all
RFMerrs.png
Attachment 2: IFOrecovery.png  283 kB  | Hide | Hide all | Show all
IFOrecovery.png
Attachment 3: ALS_ool.pdf  144 kB  | Show | Hide all | Show all
Attachment 4: PRMIcarr.png  167 kB  | Show | Hide all | Show all
ELOG V3.1.3-